what qualntity of a 65% acid solution must be mixed with a 35% solution to produce 300 mL of a 60% solution?


Answer by checkley79(3341)

.65X+.35(300-X)=300*.60
.65X+105-.35X=180
.30X=180-105
.30X=75
X=75/.30
X=250 mL OF .65% ACID IS USED.
300-250=50 mL OF .35% ACID IS USED.